# Recent Advances in OOD Detection: Problems and Approaches

### v2

@@ 2. Related work
Anomaly Detection. Anomaly detection (AD) involves identifying data points, events, or observations that deviate significantly from the dataset’s normal behavior  (Chandola et al., 2009). These anomalies can indicate critical incidents, such as fraud  (Pourhabibi et al., 2020), network intrusions  (Lazarevic et al., 2003), or system failures  (Du et al., 2017). The process involves using statistical, machine learning, or deep learning methods to model normal behavior and detect deviations. Anomaly detection is crucial in domains like cybersecurity, finance, healthcare, and manufacturing, where recognizing unusual patterns quickly can prevent significant losses or improve operational efficiency. While both AD and OOD detection aim to identify unusual or unexpected data points, AD focuses on deviations within a single distribution (typically the training set), whereas OOD detection addresses differences between the training distribution and new, unseen inputs.
@@
Open Set Recognition. Open set recognition (OSR) extends beyond traditional classification tasks by not only classifying known classes but also recognizing when input does not belong to any of the known categories (Bendale and Boult, 2016). This is crucial for real-world applications where the environment is dynamic, and the system encounters instances that were not present during training. Open set recognition is essential in fields like robotics  (Marsland et al., 2005), autonomous vehicles, and image recognition systems  (Amini et al., 2018), where encountering unknown objects or scenarios is common, and the system must be able to handle these gracefully. OSR is concerned with identifying new classes that are not present during training, which is generally achieved by dividing the categories of the same dataset into base classes and new classes, which means that new classes and base classes usually come from the same domain. In contrast, OOD detection focuses on identifying any data that is different from the training distribution, regardless of whether it belongs to a new class or a different domain altogether. Essentially, OSR is a subset of OOD detection, specifically aimed at classifying new class types within the same domain.
@@ 3.1. OOD Detection Approaches with only ID Data
Overview. Given the ID data, approaches in this section train a model on them and aim to utilize the model for detecting OOD test samples while ensuring accurate classification of the ID data. Approaches with only ID data focus specifically on mining information from ID data, without explicitly relying on other information from real-world OOD data. We further differentiate these methods into the following four categories: Reconstruction-based, Probability-based, Logits-based, OOD Synthesis, and Prototype-based. Considering real-world requirements, we also delve into a specific scenario: Long-tail ID data.
@@
Reconstruction-based. The reconstruction-based methodology offers a new research avenue by scrutinizing the discrepancy between sample representations before and after reconstruction, which relies heavily on the reconstruction performance of models. Fundamentally, the objective of the reconstruction task is to recuperate the inherent semantic content of the dataset, guided by a designated supervisory signal. This premise rests on the notion that OOD data intrinsically possess semantic characteristics that are incongruent with those of ID label information. By quantitatively assessing the extent of this semantic deviation, the model becomes equipped to discern OOD data with a high degree of precision. The approximate process of such methods is shown in Fig.  3.
@@
MoodCat  (Yang et al., 2022a) masks random portions of input images and utilizes a generative model to synthesize new images based on the classification results, implementing strong constraints during the synthesis process. Similarly,  Zhou (2022) introduces an auxiliary module to extract activations of feature vectors, aiding the model in constraining the latent reconstruction space to filter potential OOD data. Following this, MOOD  (Li et al., 2023a) and MOODv2  (Li et al., 2024a) leverage pretraining tasks based on masked image modeling, exhibiting significant advantages in learning the internal distribution of data. PRE  (Osada et al., 2023) introduces the concept of normalized flow, combined with a penalty based on typicality to constrain reconstruction errors, which discerns differences between OOD and ID data well.
Figure 3. Illustration of reconstruction-based OOD detection approaches. After extracting the feature representation $F(x)$ of the image through a neural network, it is then fed into a reconstruction network (such as VAE or DDPM) to obtain the reconstructed image representation $F^{\prime}(x)$. By comparing the dissimilarity between the original and reconstructed images, we can identify OOD data, as OOD samples typically exhibit greater dissimilarity before and after reconstruction.
@@
Recently, diffusion models have made remarkable strides in both training stability and the quality of generated images. Leveraging diffusion models to detect OOD data has become a new research direction.  Graham et al. (2023) address this issue by introducing DDPM, utilizing it for the reconstruction of noise-disturbed images. The model employs multidimensional reconstruction error for the identification of OOD data. The information bottleneck of this model can be externally regulated. Similarly, LMD  (Liu et al., 2023a) utilizes diffusion models for OOD validation. LMD disrupts data and then employs diffusion models to reconstruct images separated from the original manifold, distinguishing OOD data by comparing differences with the original manifold. DiffGuard  (Gao et al., 2023a) directly employs a pre-trained diffusion model for semantic mismatch guidance, aiming to leverage the diffusion model to amplify the disparity between reconstructed OOD images and the original images.
@@
Probability-based. Research in probability-based direction aims to establish probability models to describe the distribution of training data. Through these probability models, suitable scoring functions are developed to compute the scores of test samples within the ID distribution, which reflect whether these samples belong to the ID distribution. In this domain,  Li et al. (2023b) model each ID category during training using multiple Gaussian mixture models, and during the prediction phase, it combines Mahalanobis distance metrics to assess the likelihood of anomalous classes.  Huang et al. (2022) address this issue by introducing two regularization constraints. The density consistency regularization aligns the analytical density with low-dimensional class labels, and the contrastive distribution regularization helps separate the density between ID and OOD samples. Furthermore, LID (Kamkari et al., 2024) introduces a new detection criterion for the OOD detection paradox in the context of data generation by deep generative models. It measures whether data should be classified as in-distribution by estimating the local intrinsic dimension (LID) of the learned manifold of the generative model, when the data is assigned a high probability and the probability mass is non-negligible.
@@
Logits-based. These algorithms primarily focus on the predictions of neural networks, particularly on Logits, which are the results of the neural network’s output layer. Logits typically represent the model’s confidence or probability for each category. LogitNorm  (Wei et al., 2022a) proposes a method using logit normalization to enforce a constant vector norm on logits during training to mitigate issues of model overconfidence. Similarly, UE-NL  (Huang and Qiao, 2023), derived from Bayesian networks, normalizes logits while simultaneously learning embeddings and uncertainty scores. It adjusts the learning intensity between samples during training, rendering the learning process robust. DML  (Zhang and Xiang, 2023) challenges the potential hindrance to OOD detection performance. Drawing on empirical insights, DML enhances OOD detection performance by decoupling logits and balancing individual components, thereby mitigating the impact of each attribute on the results. It enhances OOD detection performance by decoupling logits and scenario MaxNorm as a parameter, thereby balancing the influence of each attribute on the outcome.
@@
OOD Synthesis. In the task of OOD detection, it is theoretically believed that incorporating the features of OOD data during model training can enhance the OOD detection performance. Due to the challenges in acquiring distribution information for OOD samples, some methods employ ID data to estimate the distribution of OOD data. This is done to simulate the scenario in which a model encounters OOD data in real-world situations.
Figure 4. Illustration of OOD synthesis approaches. In the absence of real-world OOD data, utilizing extrapolation from ID data to identify suitable OOD data and then training the model using the method of exposing anomalies enhances the capability of OOD detection. The orange dots in the figure represent virtual OOD data synthesized using ID data.
@@
Lee et al. (2018a) add additional regularization terms by incorporating two supplementary terms into the original loss function. The first term compels the classifier to increase uncertainty for OOD samples, while the second term encourages the GAN to generate appropriate OOD samples. Specifically, a joint training scheme is designed to select suitable boundary samples in the low-density regions of ID samples. Similarly, VOS  (Du et al., 2022a) employs adaptive synthesis of anomaly values. It samples anomaly values from the low-likelihood regions of class-conditional distributions in the feature space, eliminating the need for reliance on external data. This enables the model to synthesize OOD data for predictions. It is worth noting that VOS supports adaptive outlier synthesis, making it easily applicable to any ID data without the need for manual intervention. Noticing the additional distributional assumptions posed by previous methods for OOD data, Some researches approach the problem from another perspective. For example, NPOS  (Tao et al., 2023) contends that, rather than modeling the feature space of outliers as a parameterized Gaussian distribution, not making any distribution assumptions about ID embedded data would endow the model with robust flexibility and generality. it selects boundary points through non-parametric density estimation based on nearest neighbors, with these boundary points situated between ID and OOD data. Additionally, SSOD  (Pei, 2024) introduces self-supervised sampling to implicitly generate OOD data. It directly samples natural OOD signals from the backgrounds of ID data images, thus overcoming limitations introduced by biases during the OOD synthesis stage.
@@
In contrast to the implicit generation methods of OOD samples mentioned above, some approaches take a different route by directly leveraging explicit construction of OOD samples from ID samples for learning. To acquire OOD samples, CODEs  (Tang et al., 2021) initially generate seed OOD examples by slicing and stitching samples from internal distributions of different categories. These examples are then inputted into Chamfer GAN for distribution transformation, yielding high-quality OOD samples. CMG  (Wang et al., 2022a) then generates pseudo OOD data by providing class embeddings mixed as conditions to a Conditional Variational Autoencoder (CVAE), and subsequently utilizes this data to fine-tune the classifier for OOD detection. SEM  (Yang et al., 2023b) introduces a problem setting for full-spectrum OOD detection, where negative samples are generated using Mixup operations during training. It then leverages both high-level semantic information and low-level non-semantic features, such as style, to identify the source of the data. SHIFT  (Kwon et al., 2023) proposes a direct synthesis of OOD image samples based on training examples. It achieves this by employing CLIP to eliminate ID object regions within the training samples. The latent diffusion model is then utilized to substitute these regions with authentic features, all the while taking into account the contextual background. This methodology thereby establishes the model’s capability for rejection. However, ensuring the quality of generated data is often challenging. Furthermore, the quality of the generated data itself may have inherent flaws. ATOL  (Zheng et al., 2023) introduces an auxiliary task, wherein both auxiliary ID and auxiliary OOD data coexist. In a low-dimensional latent space, distinct regions are manually sought for both auxiliary ID and auxiliary OOD data, ensuring non-overlapping characteristics in the input space. Subsequently, data generation by the generator guarantees the non-overlapping properties of the input space. Finally, reliability is upheld by aligning genuine ID data with auxiliary ID data, effectively alleviating issues related to erroneously generated OOD instances.
@@
Prototype-based. During the model training process, prototype-based OOD detection methods aim to model the ID data using prototypes to learn the common distribution characteristics of the ID data. In the testing phase, the model measures the differences between the sample and class-level prototypes to determine the category of the sample. The general procedure of these methods is roughly illustrated in Fig.  5.
Figure 5. Illustration of prototype-based OOD detection approaches. The algorithm for OOD detection based on prototype learning roughly consists of three network structures: 1) Encoder. This structure is responsible for propagating sample data through a neural network to obtain feature representations $h$. 2) Projector. Structures of this type project the image features encoded in $h$ to a new representation space. 3) Prototype Learning. In the new representation space, prototypes representing each class are learned, with each class learning its corresponding prototype center. During testing, the likelihood of anomalous data is measured by comparing the similarity between test samples and prototypes of each ID class.
@@
SIREN  (Du et al., 2022b) first models the distribution of ID data using the von Mises-Fisher(vMF) model, which enables representing each class as a compact cluster, i.e., a class prototype. Generally, the vMF distribution modeling formula can be represented as: $p_{D}(z;p_{k},k)=Z_{D}(k)exp(kp_{k}^{T}z)$, where $p_{k}$ is the $k$-th prototype with unit norm, $\kappa\geq 0$ represents the concentration around the mean, and $Z_{D}(\kappa)$ is the normalization factor. In the prototype-based approach, an embedding vector $z$ is assigned to class $c$ with the following normalized probability:
@@
where $c\in{\{1,2,...C\}}$. Additionally, within the loss function, it enforces alignment between the embedding vectors of ID samples and class prototypes, constraining each ID class sample. This parametric OOD score can be directly obtained after training, without requiring separate estimation. CIDER  (Ming et al., 2023) builds upon SIREN by jointly optimizing two losses to enhance data discriminability, encouraging the maximization of angular distances between prototypes of different classes and the internal compactness of prototypes within the same class. The optimization process during model training pertains to the prototypes of classes. ReweightOOD  (Regmi et al., 2024) argues that optimizing non-class data impedes achieving clear class separability, while focusing on fewer class data makes it challenging to achieve lower MSE scores. To address this, they propose a re-weighting optimization strategy to balance the significance of different losses. Although the ideas behind Step  (Zhou et al., 2021) are different, in the context of semi-supervised tasks, it essentially generates clusters of unlabeled ID and OOD samples through a contrastive learning process, which is conceptually similar to prototype learning.
However, PALM  (Lu et al., 2024) observes that modeling each class with a single prototype may fail to capture the diversity of internal information within the data. PALM introduces a mixed-prototype strategy to optimize the modeling process, utilizing a strategy of multiple prototypes to learn informative representations for each class. By simultaneously learning class-level prototypes and contrasting inter-class prototypes, it optimizes prototype-level intra-class compactness and inter-class discriminability within the loss function.

### v3

@@ 2. Related work
Anomaly Detection. Anomaly detection (AD) involves identifying data points, events, or observations that deviate significantly from the dataset’s normal behavior  (Chandola et al., 2009; Xing et al., 2025; Wang and Nalisnick, 2025). These anomalies can indicate critical incidents, such as fraud  (Pourhabibi et al., 2020), network intrusions  (Lazarevic et al., 2003) or system failures  (Du et al., 2017). The process involves using statistical, machine learning, or deep learning methods to model normal behavior and detect deviations. Anomaly detection is crucial in domains like cybersecurity, finance, healthcare, and manufacturing, where recognizing unusual patterns quickly can prevent significant losses or improve operational efficiency. While both AD and OOD detection aim to identify unusual or unexpected data points, compared to OOD detection, AD differs in three key aspects. First, regarding supervision, AD methods are typically unsupervised, training only on normal samples without detailed labels and detecting anomalies during testing, whereas OOD detection involves labeled category discrimination. Second, AD focuses on perceptual anomalies, where images belong to known categories but contain anomalous regions, while OOD detection targets semantic differences, with test samples from unseen categories. Additionally, although both methods commonly use AUROC (Wehler et al., 2001) for evaluation, AD specifically employs AUPRO (Bertoldo et al., 2024) to assess pixel-level detection performance.
@@
Open Set Recognition. Open set recognition (OSR) extends beyond traditional classification tasks by not only classifying known classes but also recognizing when input does not belong to any of the known categories (Bendale and Boult, 2016). This is crucial for real-world applications where the environment is dynamic, and the system encounters instances that are not present during training. Open set recognition is essential in fields like robotics  (Marsland et al., 2005), autonomous vehicles, and image recognition systems  (Amini et al., 2018), where encountering unknown objects or scenarios is common, and the system must be able to handle these gracefully. OSR is concerned with identifying new classes that are not present during training, which is generally achieved by dividing the categories of the same dataset into base classes and new classes, which means that new classes and base classes usually come from the same domain. In contrast, OOD detection focuses on identifying any data that is different from the training distribution, regardless of whether it belongs to a new class or a different domain altogether. Essentially, OSR is a subset of OOD detection, specifically aimed at classifying new class types within the same domain.
@@
Misclassification (Failure) Detection. Misclassification detection aims to identify incorrect predictions on in-distribution data, exposing a model’s unreliability in its intended domain (Ahmed et al., 2022; Dong et al., 2025; Inceoglu et al., 2021; Liu et al., 2024d). This is especially important in safety-critical areas such as autonomous driving and healthcare. Failure detection generalizes this task to capture a broader range of model breakdowns, including uncertainty and data corruption. While both misclassification and OOD detection seek to improve model reliability, the former focuses on errors within the training distribution (Jaeger et al., 2023), whereas the latter targets inputs outside it.
@@ 3.1. OOD Detection Approaches with only ID Data
Overview. Given the ID data, approaches in this section train a model on them and aim to utilize the model for detecting OOD test samples while ensuring classification performance of the ID data. They focus specifically on mining information from ID data, without explicitly relying on other information from real OOD data. We further differentiate these methods into the following five categories: Reconstruction-based, Probability-based, Logits-based, OOD Synthesis, and Prototype-based. Considering real-world requirements, we also delve into a specific scenario: Long-tail ID data.
@@
Reconstruction-based. The reconstruction-based approach detects OOD samples by measuring the difference between representations of data before and after reconstruction, relying on the model’s ability to accurately recover input semantics. The core idea is that OOD data contain semantic features that differ from those of ID data. By quantifying these differences, the model can effectively distinguish OOD samples. An overview is illustrated in Fig. 3 (a).
@@
MoodCat  (Yang et al., 2022) masks random portions of input images and utilizes a generative model to synthesize new images based on the classification results, implementing strong constraints during the synthesis process. Similarly,  Zhou (2022) introduces an auxiliary module to extract activations of feature vectors, aiding the model in constraining the latent reconstruction space to filter potential OOD data. Following this, MOOD  (Li et al., 2023a) and MOODv2  (Li et al., 2024a) leverage pretraining tasks based on masked image modeling, exhibiting significant advantages in learning the internal distribution of data. PRE  (Osada et al., 2023) introduces the concept of normalized flow, combined with a penalty based on typicality to constrain reconstruction errors, which discerns differences between OOD and ID data well.
Figure 3. Illustration of some subclasses within the Training-driven approaches: (a) Reconstruction-based : After extracting the feature representation $F(x)$ of the image, it is fed into a reconstruction network (e.g., VAE or DDPM) to obtain the reconstructed feature representation $F^{\prime}(x)$. By comparing the dissimilarity between the original and reconstructed images, OOD data can be identified, as OOD samples typically exhibit higher dissimilarity before and after reconstruction. (b) Prototype-based : The model learns more optimal ID prototypes during the training process to distinguish OOD data.The symbol $h$ typically denotes the features of the penultimate layer in the network. (c) OOD Synthesis : During training, real OOD data is not available; instead, OOD data is synthesized by extrapolating information from the ID data. (d) Outlier Mining : During training, real OOD data is available, but typical OOD samples are generated using this real OOD data. (e) Boundary Regularization : The model’s decision boundary is optimized by using real OOD data, typically without involving the synthesis of OOD data.
@@
Recently, diffusion models (Ho et al., 2020) have made remarkable strides in both training stability and the quality of generated images. Leveraging diffusion models to detect OOD data has become a new research direction.  Graham et al. (2023) address this issue by introducing DDPM, utilizing it for the reconstruction of noise-disturbed images. The model employs multidimensional reconstruction error for the identification of OOD data. The information bottleneck of this model can be externally regulated. Similarly, LMD  (Liu et al., 2023a) utilizes diffusion models for OOD validation. LMD disrupts data and then employs diffusion models to reconstruct images separated from the original manifold, distinguishing OOD data by comparing differences with the original manifold. DiffGuard  (Gao et al., 2023b) directly employs a pre-trained diffusion model for semantic mismatch guidance, aiming to leverage the diffusion model to amplify the disparity between reconstructed OOD images and the original images.
@@
Probability-based. Probability-based approaches aim to establish probability models to describe the distribution of training data. In this domain,  Li et al. (2023b) model each ID category during training using multiple Gaussian mixture models (Rasmussen, 1999), and during the prediction phase, it combines Mahalanobis distance metrics to assess the likelihood of anomalous classes.  Huang et al. (2022a) address this issue by introducing two regularization constraints. The density consistency regularization aligns the analytical density with low-dimensional class labels, and the contrastive distribution regularization helps separate the density between ID and OOD samples. Furthermore, LID (Kamkari et al., 2024) introduces a new detection criterion for the OOD detection paradox in the context of data generation by deep generative models. It measures whether data should be classified as in-distribution by estimating the local intrinsic dimension (LID) of the learned manifold of the generative model, when the data is assigned a high probability and the probability mass is non-negligible.
@@
Logits-based. These methods detect OOD samples by analyzing the logits (raw outputs before softmax) from the neural network’s final layer. Logits typically represent the model’s confidence or probability for each category. LogitNorm  (Wei et al., 2022a) uses logit normalization to enforce a constant vector norm on logits during training to mitigate issues of model overconfidence. Similarly, UE-NL  (Huang and Qiao, 2023), derived from Bayesian networks, normalizes logits while simultaneously learning embeddings and uncertainty scores. It adjusts the learning intensity between samples during training, rendering the learning process robust. DML  (Zhang and Xiang, 2023) challenges the potential hindrance to OOD detection performance. Drawing on empirical insights, DML enhances OOD detection performance by decoupling logits and balancing individual components, thereby mitigating the impact of each attribute on the results. It enhances OOD detection performance by decoupling logits and scenario MaxNorm as a parameter, thereby balancing the influence of each attribute on the outcome.
@@
OOD Synthesis. In the task of OOD detection, incorporating the features of OOD data during model training can enhance the OOD detection performance by allowing the model to better recognize and differentiate OOD data. Due to the challenges in acquiring distribution information for OOD samples, some methods employ ID data to estimate the distribution of OOD data, simulating real-world scenarios where a model encounters OOD data, as shown in Fig. 3 (c).
@@
Lee et al. (2018a) add additional regularization terms by incorporating two supplementary terms into the original loss function. The first term compels the classifier to increase uncertainty for OOD samples, while the second term encourages the GAN to generate appropriate OOD samples. Specifically, a joint training scheme is designed to select suitable boundary samples in the low-density regions of ID samples. Similarly, VOS  (Du et al., 2022a) employs adaptive synthesis of anomaly values. It samples anomaly values from the low-likelihood regions of class-conditional distributions in the feature space, eliminating the need for reliance on external data. This enables the model to synthesize OOD data for predictions. It is worth noting that VOS supports adaptive outlier synthesis, making it easily applicable to any ID data without the need for manual intervention. Noticing the additional distributional assumptions posed by previous methods for OOD data, some researchers approach the problem from another perspective. For example, NPOS  (Tao et al., 2023) contends that, rather than modeling the feature space of outliers as a parameterized Gaussian distribution, not making any distribution assumptions about ID embedded data would endow the model with robust flexibility and generality. It selects boundary points through non-parametric density estimation based on nearest neighbors, with these boundary points situated between ID and OOD data. Additionally, SSOD  (Pei, 2024) introduces self-supervised sampling to implicitly generate OOD data. It directly samples natural OOD signals from the backgrounds of ID data images, thus overcoming limitations introduced by biases during the OOD synthesis stage.
@@
In contrast to the implicit generation methods of OOD samples mentioned above, some approaches take a different route by directly leveraging explicit construction of OOD samples from ID samples for learning. To acquire OOD samples, CODEs  (Tang et al., 2021) initially generate seed OOD examples by slicing and stitching samples from internal distributions of different categories. These examples are then inputted into Chamfer GAN for distribution transformation, yielding high-quality OOD samples. CMG  (Wang et al., 2022a) then generates pseudo OOD data by providing class embeddings mixed as conditions to a Conditional Variational Autoencoder (CVAE), and subsequently utilizes this data to fine-tune the classifier for OOD detection. SEM  (Yang et al., 2023b) introduces a problem setting for full-spectrum OOD detection, where negative samples are generated using Mixup operations during training. It then leverages both high-level semantic information and low-level non-semantic features, such as style, to identify the source of the data. Forte (Ganguly et al., 2025) utilizes per-point statistics (such as precision, recall, and density) to capture the local geometric properties of the data manifold in the feature space and performs typicality estimation using density-based non-parametric methods (such as Gaussian mixture models). This approach helps improve the accuracy of OOD synthetic data detection. SHIFT  (Kwon et al., 2023) proposes a direct synthesis of OOD image samples based on training examples. It achieves this by employing CLIP to eliminate ID object regions within the training samples. The latent diffusion model is then utilized to substitute these regions with authentic features, all the while taking into account the contextual background. This methodology thereby establishes the model’s capability for rejection. However, ensuring the quality of generated data is often challenging. Furthermore, the quality of the generated data itself may have inherent flaws. ATOL  (Zheng et al., 2023) introduces an auxiliary task, wherein both auxiliary ID and auxiliary OOD data coexist. In a low-dimensional latent space, distinct regions are manually sought for both auxiliary ID and auxiliary OOD data, ensuring non-overlapping characteristics in the input space. Subsequently, data generation by the generator guarantees the non-overlapping properties of the input space. Finally, reliability is upheld by aligning genuine ID data with auxiliary ID data, effectively alleviating issues related to erroneously generated OOD instances. In recent studies, HamOS (Li and Zhang, 2025) employs Hamiltonian Monte Carlo (Duane et al., 1987) sampling, treating the synthesis process as sampling from a Markov chain. In the absence of OOD data, it synthesizes high-quality OOD virtual anomaly points by constructing a Markov chain (Neal, 1993). It projects the features onto a hyperspherical feature space and uses K-nearest neighbor distances to guide the generation of anomaly points. POP (Gong et al., 2025) does not require the synthesis of specific outliers, but instead directly reshapes the decision boundaries through virtual OOD prototypes, improving the model’s OOD detection capability by penalizing misclassifications.
@@
Prototype-based. During the model training process, prototype-based OOD detection methods aim to model the ID data using prototypes to learn the common distribution characteristics of the ID data. For OOD data, there is a significant difference between the sample features and the ID prototype. In the testing phase, the model determines the sample’s category by measuring the difference between the sample and the ID prototype. The general procedure of these methods is outlined in Fig. 3 (b).
@@
SIREN  (Du et al., 2022b) first models the distribution of ID data using the von Mises-Fisher(vMF) (Mardia and Jupp, 2009) model, which enables representing each class as a compact cluster, i.e., a class prototype. Generally, the vMF distribution modeling formula can be represented as: $p_{D}(z;p_{k},k)=Z_{D}(k)exp(kp_{k}^{T}z)$, where $p_{k}$ is the $k$-th prototype with unit norm, $\kappa\geq 0$ represents the concentration around the mean, and $Z_{D}(\kappa)$ is the normalization factor. In the prototype-based approach, an embedding vector $z$ is assigned to class $c$ with the following normalized probability:
@@
where $c\in{\{1,2,...C\}}$. Additionally, within the loss function, it enforces alignment between the embedding vectors of ID samples and class prototypes, constraining each ID class sample. This parametric OOD score can be directly obtained after training, without requiring separate estimation. CIDER  (Ming et al., 2023) builds upon SIREN by jointly optimizing two losses to enhance data discriminability, encouraging the maximization of angular distances between prototypes of different classes and the internal compactness of prototypes within the same class. The optimization process during model training pertains to the prototypes of classes. AROS (Mirzaei and Mathis, 2025) focuses more on the feature separation between ID and OOD data. It proposes using the Lyapunov stability theorem to constrain the stability of the neural network’s embedding space, ensuring that the embeddings of ID and OOD data converge to distinct stable equilibrium points. To obtain OOD data, it also employs a technique of sampling from the low-probability regions of the ID data to generate pseudo-OOD. CNC (Harun et al., 2025) proposed controlling the degree of neural collapse (NC) at different layers of the DNN network. In the Encoder, entropy regularization is introduced to mitigate NC, while in the Projector, the projection head is used to enhance the NC phenomenon, forming more compact prototype and thereby improving OOD detection performance. ReweightOOD  (Regmi et al., 2024) argues that optimizing non-class data impedes achieving clear class separability, while focusing on fewer class data makes it challenging to achieve lower MSE scores. To address this, they propose a re-weighting optimization strategy to balance the significance of different losses. Although the ideas behind Step  (Zhou et al., 2021) are different, in the context of semi-supervised tasks, it essentially generates clusters of unlabeled ID and OOD samples through a contrastive learning process, which is conceptually similar to prototype learning.
However, PALM (Lu et al., 2024) observes that modeling each class with a single prototype may fail to capture the diversity of internal information within the data. PALM introduces a mixed-prototype strategy to optimize the modeling process, utilizing a strategy of multiple prototypes to learn informative representations for each class. By simultaneously learning class-level prototypes and contrasting inter-class prototypes, it optimizes prototype-level intra-class compactness and inter-class discriminability within the loss function. PFS (Wu et al., 2025) also introduces a clustering loss that encourages features of the same ID class to cluster more closely, thereby improving the consistency of ID data representation.